# K theoretic virtual fundamental cycle of global Kuranishi chart

### v1

@@ 3.3.1. Statement of results
Let $i:\mathcal{Z}\to\mathcal{Y}$ be a closed suborbifold of $\mathcal{Y}$, whose boundary is contained in that of $\mathcal{Y}$. Let $\pi:\mathcal{N}\to\mathcal{Z}$ be a closed tubular neighborhood with $\pi^{-1}(\partial\mathcal{Z})\subseteq\partial\mathcal{Y}$. Define various projections by the following diagram.
Recall that $i_{\ast}^{fake}(H,\phi,T)=(H,\phi\circ i^{\ast},T)$.
Given an inclusion $j:\mathcal{N}\hookrightarrow\mathcal{Y}$, let $\mathcal{I}$ be the ideal of functions in $\mathbb{C}(\mathcal{Y})$ that vanish on $\mathcal{N}$. Given an element $(H,\phi,T)\in K_{0}^{fake}(\mathcal{Y})$, let $H_{\mathcal{N}}=H/\mathcal{I}H$, $\phi_{\mathcal{N}}(f)$ the action of an arbitrary extension of $f$ on the quotient space $H_{\mathcal{N}}$ via $\phi$, and $T_{\mathcal{N}}:H_{\mathcal{N}}\to H_{\mathcal{N}}$ the quotient map of $T$. We define
$\begin{array}[]{llll}j^{!}:&K_{0}^{fake}(\mathcal{Y})&\to&K_{0}^{fake}(\mathcal{N})\\ &(H,\phi,T)&\to&(H_{\mathcal{N}},\phi_{\mathcal{N}},T_{\mathcal{N}})\end{array}$
and $i^{fake,!}$ as the composition $\pi_{\ast}j^{!}$.
We also recall the wrong-way $K$-cohomological push-forward
defined in [4]. Let $\mathcal{N}_{\mathbb{C}}=B(\pi_{\mathcal{Z}}^{\ast}\mathcal{N}\otimes\mathbb{C})$, which is identified with $B^{\ast}\mathcal{N}\to B^{\ast}\mathcal{Z}$ by identifying the real part with fibers of $\mathcal{N}\to\mathcal{Z}$ and imaginary part with corresponding cotangent vectors. We define the K-theoretical push-forward in the special case of $\mathcal{Y}=\mathcal{N}$ as
$i_{!}^{fake}:\alpha\to(\pi^{\ast}\alpha)\otimes(str_{1}\wedge^{\ast}\mathcal{N}_{\mathbb{C}}^{\ast})$
The push-forward $i_{!}^{fake}$ in the general case is the composition of $i_{!}^{fake}$ for tubular neighborhood and the map $K^{0,fake}(B^{\ast}\mathcal{N},\partial B^{\ast}\mathcal{N})\to K^{0,fake}(B^{\ast}\mathcal{Y},\partial B^{\ast}\mathcal{Y})$ given by extending a virtual bundle trivially outside $\mathcal{N}$. $i_{!}^{fake}$ is independent of the tubular neighborhood chosen [3].
We define the genuine $i_{!},i_{\ast},i^{!}$ sector-wise, and will show in this subsection that
@@
    The homological $i^{!}$ is well defined, and the following diagram commutes.
@@ 3.3.2. A technical lemma
We first fix some notations. Fix a symbolic representative $(\pi_{\mathcal{Z}}^{\ast}\mathcal{E},\alpha)\in K^{0,fake}(B^{\ast}\mathcal{Z},\partial B^{\ast}\mathcal{Z})$. We define $\beta$ and $\tilde{D}$, and a specific $\tilde{T}$ (and $T$) that fit into the following diagram. Note that the ‘product’ $\otimes^{\prime}$ is as defined in Section 3.1 and is not the genuine tensor product, as $\tilde{T}$ and $\tilde{D}$ are not $\mathbb{C}(\mathcal{N})$-linear. Nonetheless, $\otimes^{\prime}$ is the genuine tensor product $\otimes$ after pushing-forward to $K_{0}^{fake}(\mathcal{Z})$ under $\pi_{\ast}$.
${(\pi_{\mathcal{Z}}^{\ast}\mathcal{E},\alpha)}$${(\tilde{\pi}^{\ast}\pi^{\ast}_{\mathcal{Z}}\mathcal{E}\otimes str_{1}\wedge^{\ast}\mathcal{N}_{\mathbb{C}}^{\ast},\alpha\otimes 1-\varepsilon 1\otimes\beta)}$${(L^{2}(\mathcal{Z},\mathcal{E}),T)}$${(L^{2}(\mathcal{N},\pi^{\ast}(\mathcal{E}\otimes str_{1}(\wedge^{\ast}\mathcal{N}^{\ast}\otimes\mathbb{C}))),\tilde{T}\otimes^{\prime}1-\varepsilon 1\otimes^{\prime}\tilde{D})}$$\scriptstyle{i_{!}^{fake}}$$\scriptstyle{\mu_{\mathcal{Z}}}$$\scriptstyle{\mu_{\mathcal{N}}}$
Let $\beta$ be a symbolic representative of $\lambda_{\xi}:\wedge^{\ast}\mathcal{N}_{\mathbb{C}}\to\wedge^{\ast}\mathcal{N}_{\mathbb{C}}$ (as defined in the introduction) in $K^{0,fake}(\mathcal{N})$ and $\tilde{D}$ its quantization. Then $\tilde{D}$ is a self-adjoint 0-th order pseudo-differential operator over $\mathcal{N}$ that acts in the fiber direction of $\mathcal{N}\to\mathcal{Z}$.
We write $\mathcal{Z}=Z/G,\mathcal{N}=N/G$ and pick an equivariant, order 00 pseudo-differential operator $P$ on $Z$ whose symbol represents $\alpha$. Recall that in Section 2.5, we defined $T$ as (the average of) $P$ restricted to $G$-invariant $L^{2}$-sections. To define $\tilde{T}$, we pick $\mathcal{N}^{\prime}$ such that $\mathcal{N}\oplus\mathcal{N}^{\prime}=Z\times_{G}V$ for $V$ (the unit ball in) some $G$-representation. We construct $\tilde{P}$ as the order 00 pseudo-differential operator on $Z\times V$ given by $P$ on each slice $Z\times\{v\}$, and $\tilde{P}^{G}$ as the average of $\tilde{P}$ over $G$. We define $\tilde{T}$ by its action on a section of $\pi^{\ast}\mathcal{E}$ over $\mathcal{N}$ equals the restriction to $\mathcal{N}$ of $\tilde{P}_{G}$ acting on the pull-back of that section along $\mathcal{N}\oplus\mathcal{N}^{\prime}\to\mathcal{N}$.
@@
For an element $(\mathcal{H},\phi,\mathcal{T})\in K^{fake}_{0}(\mathcal{Z})$, let $\Delta\in L(\mathcal{H})$ be a degree $1$ self adjoint operator that commutes with $\phi$ and $\mathcal{T}$. Then $(\mathcal{H},\phi,\mathcal{T}-\varepsilon\Delta)=(\ker\Delta,\phi|_{\ker\Delta},\mathcal{T})$.
@@
The operators $\mathcal{T}$ and $\Delta$ commute, so $\ker\Delta$ and $\operatorname{im}\Delta$ are both invariant under $\mathcal{T}-\varepsilon\Delta$ and $\phi(f)$. Thus $(\ker\Delta,\phi|_{\ker\Delta},\mathcal{T})$ and $(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},\mathcal{T}-\varepsilon\Delta)$ both represent elements in $K_{0}^{fake}(\mathcal{Z})$, and
@@
Now $(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},\mathcal{T}-\varepsilon\Delta)$ is homotopic to the degenerated element $(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},-\varepsilon\Delta)$ by the line segment connecting them, so
@@
Observe that the image under $\pi_{\ast}$ of $1\otimes\tilde{D}$ commutes with the images of $\tilde{T}\otimes 1$ and $\phi$, so by the previous lemma we know that $\pi_{\ast}\mu_{\mathcal{N}}i_{!}^{fake}(\mathcal{E},\alpha)=(\ker 1\otimes\tilde{D},\tilde{T}\otimes 1|_{\ker 1\otimes\tilde{D}})$.
Let $s_{i}$’s be a $\mathbb{C}(\mathcal{Z})$-basis of $C(\mathcal{N})$, and form $\beta_{i},\tilde{D}_{i}$ from $s_{i}$’s as we form $\beta,\tilde{D}$. Then $1\otimes\tilde{D}_{i}$ commutes with $\tilde{T}\otimes 1$ and $\phi$. If $(V,\tilde{T}\otimes 1)\in K_{0}^{fake}(\mathcal{Z})$, then the self-adjoint operator $\tilde{T}\otimes 1$ almost commutes with the $\mathbb{C}(\mathcal{Z})$-action, and so represents the same element as $(V,\tilde{T}\otimes 1-\varepsilon\tilde{D}_{i})$. Thus, again, by the previous lemma, we can restrict $V$ to $\ker\tilde{D_{i}}$. So $\pi_{\ast}\mu_{\mathcal{N}}i_{!}^{fake}(\mathcal{E},\alpha)=(\cap_{i}\ker(1\otimes\tilde{D}_{i}),\tilde{T}\otimes 1)$.
On each fiber of $\mathcal{N}\to\mathcal{Z}$, the equations given by $\tilde{D}_{i}=0$ precisely determine a section over $\mathcal{N}$ from its initial values in $\mathcal{Z}$ ([4]). Thus the common kernels of $\tilde{D}_{i}$ is isomorphic to $(L^{2}(\mathcal{Z},\mathcal{E}),\tilde{T}\otimes 1)=(L^{2}(\mathcal{E}),T\otimes 1)$, and the proposition follows. ∎
@@
It suffices to prove the statement sector-wise, and that the statement on twisted sectors is identical to that on the non-twisted sector (only with $\mathcal{Z},\mathcal{N},\mathcal{Y}$’s replaced by the corresponding twisted sectors in these spaces).
@@
Observe that the contraction $\pi:\mathcal{N}\to\mathcal{Z}$ composed with $i^{fake}:\mathcal{Z}\hookrightarrow\mathcal{N}$ is homotopic to $id:\mathcal{N}\to\mathcal{N}$, thus induces a homotopy $\pi^{\ast}(i^{fake})^{\ast}\cong id$ on $\mathbb{C}(\mathcal{N})$. By [10] Theorem 3 in Section 3, we know that $i^{fake}_{\ast}\pi_{\ast}=id$ on $K_{0}^{fake}(\mathcal{N})$. On the other hand, it follows from $\pi i^{fake}=id_{\mathcal{Z}}$ that $\pi_{\ast}i^{fake}_{\ast}=id$ on $K_{0}^{fake}(\mathcal{Z})$. So $\pi_{\ast}$ and $i^{fake}_{\ast}$ are inverse to each other, and the corollary follows. ∎
@@
Part (1) follows from the definition. If furthermore $\mathcal{Y}$ deformation retracts to $\mathcal{N}$, then $j_{\ast}$ is an isomorphism by [10] Theorem 3 in Section 3. Thus $j_{\ast}=(j^{!})^{-1}$ by part (1). ∎
Proof of $i^{fake,!}$ being independent of the choice of $\mathcal{N}$.
Different tubular neighborhood structures on the same set $\mathcal{N}$ define the same map $i^{fake,!}$. The map $i^{fake,!}$ defined by $\mathcal{N}^{\prime}\subseteq\mathcal{N}$ and that defined by $\mathcal{N}$ coincide, by the second part of the previous lemma. Next, for two general tubular neighborhoods given, we can always find a tubular neighborhood contained in both. ∎
@@
The following diagrams commutes for each open embedding $\mathcal{N}\hookrightarrow\mathcal{Y}$.
@@
The commutativity of the first diagram follows Corollary 2.2.
The second diagram’s commutativity follows that of the first diagram and the fact that $j!j_{a}st$ is the identity map. ∎
The well-definedness part of Proposition 3.3 follows from Lemma 3.5, Corollary 3.1, and Lemma 3.8.
@@ 3.4. Push-forward from inertia orbifold to point
Consider sequences $(g,g_{1}G_{1},\cdots,g_{n}G_{n})$ of cosets in $G$, where $G_{i+1}$ is the group generated by the previous term $g_{i}G_{i}$, and $G_{1}$ the group generated by $g$. Let $\mathcal{G}$ be such sequences with $g_{i}\notin G_{i}$. Given a sequence $\mathbf{g}=(g,g_{1},\cdots,g_{n})$, we can construct a sequence of such cosets by setting $G_{i}$ the group generated by $g,g_{1},\cdots,g_{i}$. The sequence of elements $\mathbf{g}$ defines an element in $\mathcal{G}$, if and only if each $g_{i}$ is not in the group generated by previous elements.
Given $\mathbf{g}$, let $\mathcal{Y}^{\mathbf{g}}$ be the $(n+1)$-multi-sector corresponding to $\mathbf{g}$, and $i_{\mathbf{g}}:\mathcal{Y}^{\mathbf{g}}\to\mathcal{Y}$. Let $n=|\mathbf{g}|$ and $m(\mathbf{g})$ be the order of isotropy of $\mathcal{Y}^{\mathbf{g}}$. Let $\mathcal{N}_{g_{i}}$ be the normal bundle of $\mathcal{Y}^{(g,g_{1},\cdots,g_{i})}$ in $\mathcal{Y}^{(g=g_{0},g_{1},\cdots,g_{i-1})}$, and $\mathcal{N}_{g}$ in particular be the normal bundle of $\mathcal{Y}^{(g)}$ in $\mathcal{Y}$. Define $e^{-1}_{\mathbf{g}}=(\prod_{i=0}^{n}str_{g_{i}}\wedge^{\ast}(\mathcal{N}^{\ast}_{g_{i}}\otimes\mathbb{C}))^{-1}$. These objects only depend on the element in $\mathcal{G}$ represented by $\mathbf{g}$, and are independent of the specific representative $\mathbf{g}$. We define the fake push-forward $p_{\ast}^{fake}$ of an element $(H,\phi,T)$ in the fake $K$-homology as the Fredholm index of $T_{+}:H_{+}\to H_{-}$, where $H_{\pm}$ are the positive and negative part of $H$ under the $\mathbb{Z}_{2}$ grading, and $T_{+}=T|_{H_{+}}$.
@@
The orbifold ‘push-forward to point’ $p_{\ast}(\varphi)$ of $\{\varphi^{(g)}=(H^{(g)},\phi^{(g)},T^{(g)})\}\in K_{0}(\mathcal{Y})$ is
$p_{\ast}(\varphi)=p_{\ast}^{fake}\left(\sum_{\mathbf{g}\in\mathcal{G}}\frac{(-1)^{|\mathbf{g}|+\dim\mathcal{Y}^{(g)}}}{m(\mathbf{g})}(e_{\mathbf{g}}^{-1}i_{\mathbf{g}}^{!}\varphi^{(g)})\right).$
@@
The following diagram commutes, where $\mathbb{Z}\to\mathbb{C}$ is the natural inclusion, and $\operatorname{ind}(\cdots)_{+}$ maps $u$ to the index of $u_{+}$.

### v2

@@ 3.4.1. Statement of results
We will prove the following proposition in this subsection.
@@ 3.4.2. A technical lemma
Let $\pi:\mathcal{N}\to\mathcal{Z}$ be a closed tubular neighborhood of $\mathcal{Z}$. Define various projections by the following diagram.
@@
We remark that composing with the index map $K^{fake}_{0}(\mathcal{X})\to\mathbb{C}$ given by the (virtual) Fredholm index of $T_{+}:H_{+}\to H_{-}$ we know that $i_{!}^{fake}$ preserves indices. Thus this lemma could be viewed as a strengthening of corresponding theorems in index theory (e.g., Theorem 4.3 in [3] restricted to orbifolds).
We first fix some notations. Fix a symbolic representative
$(\pi_{\mathcal{Z}}^{\ast}\mathcal{E},\alpha)\in K^{0,fake}(B^{\ast}\mathcal{Z},\partial B^{\ast}\mathcal{Z}).$
We define $\beta$ and $\tilde{D}$, and a specific $\tilde{T}$ (and $T$) that fit into the following diagram. Note that the ‘product’ $\otimes^{\prime}$ is as defined in Section 3.1 and is not the genuine tensor product, as $\tilde{T}$ and $\tilde{D}$ are not $\mathbb{C}(\mathcal{N})$-linear. Nonetheless, $\otimes^{\prime}$ is the genuine tensor product $\otimes$ after pushing forward to $K_{0}^{fake}(\mathcal{Z})$ under $\pi_{\ast}$.
${(\pi_{\mathcal{Z}}^{\ast}\mathcal{E},\alpha)}$${(\tilde{\pi}^{\ast}\pi^{\ast}_{\mathcal{Z}}\mathcal{E}\otimes str_{1}\wedge^{\ast}\mathcal{N}_{\mathbb{C}}^{\ast},\alpha\otimes 1-\varepsilon 1\otimes\beta)}$${(L^{2}(\mathcal{Z},\mathcal{E}),T)}$${\Bigl{(}L^{2}\bigl{(}\mathcal{N},\pi^{\ast}(\mathcal{E}\otimes str_{1}(\wedge^{\ast}\mathcal{N}^{\ast}\otimes\mathbb{C}))\bigr{)},\tilde{T}\otimes^{\prime}1-\varepsilon 1\otimes^{\prime}\tilde{D}\Bigr{)}}$$\scriptstyle{i_{!}^{fake}}$$\scriptstyle{\mu_{\mathcal{Z}}}$$\scriptstyle{\mu_{\mathcal{N}}}$
Let $\beta$ be a symbolic representative of $\lambda_{\xi}:\wedge^{\ast}\mathcal{N}_{\mathbb{C}}\to\wedge^{\ast}\mathcal{N}_{\mathbb{C}}$ (as defined in the introduction) in $K^{0,fake}(\mathcal{N})$ and let $\tilde{D}$ be its quantization. Then $\tilde{D}$ is a self-adjoint 0th order pseudo-differential operator over $\mathcal{N}$ that acts in the fiber direction of $\mathcal{N}\to\mathcal{Z}$.
We write $\mathcal{Z}=Z/G$ and $\mathcal{N}=N/G$ and pick an equivariant, order 00 pseudo-differential operator $P$ on $Z$ whose symbol represents $\alpha$. Recall that in Section 2.5, we defined $T$ as (the average of) $P$ restricted to $G$-invariant $L^{2}$-sections. To define $\tilde{T}$, we pick $\mathcal{N}^{\prime}$ such that $\mathcal{N}\oplus\mathcal{N}^{\prime}=Z\times_{G}V$ for $V$ (the unit ball in) some $G$-representation. We construct $\tilde{P}$ as the order 00 pseudo-differential operator on $Z\times V$ given by $P$ on each slice $Z\times\{v\}$, and $\tilde{P}^{G}$ as the average of $\tilde{P}$ over $G$. We define $\tilde{T}$ by requiring that its action on a section of $\pi^{\ast}\mathcal{E}$ over $\mathcal{N}$ equals the restriction to $\mathcal{N}$ of $\tilde{P}_{G}$ acting on the pull-back of that section along $\mathcal{N}\oplus\mathcal{N}^{\prime}\to\mathcal{N}$.
@@
For an element $(\mathcal{H},\phi,\mathcal{T})\in K^{fake}_{0}(\mathcal{Z})$, let $\Delta\in L(\mathcal{H})$ be a degree $1$ self-adjoint operator that commutes with $\phi$ and $\mathcal{T}$. Then
$(\mathcal{H},\phi,\mathcal{T}-\varepsilon\Delta)=(\ker\Delta,\phi|_{\ker\Delta},\mathcal{T}).$
@@
The operators $\mathcal{T}$ and $\Delta$ commute, so $\ker\Delta$ and $\operatorname{im}\Delta$ are both invariant under $\mathcal{T}-\varepsilon\Delta$ and under $\phi(f)$. Thus, $(\ker\Delta,\phi|_{\ker\Delta},\mathcal{T})$ and $(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},\mathcal{T}-\varepsilon\Delta)$ both represent elements in $K_{0}^{fake}(\mathcal{Z})$, and
@@
Now, $(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},\mathcal{T}-\varepsilon\Delta)$ is homotopic to the degenerate element $(\operatorname{im}\Delta,\phi|_{\operatorname{im}\Delta},-\varepsilon\Delta)$ by the line segment connecting them, so
@@
Observe that the image under $\pi_{\ast}$ of $1\otimes\tilde{D}$ commutes with the images of $\tilde{T}\otimes 1$ and $\phi$, so by the previous lemma we know that
$\pi_{\ast}\mu_{\mathcal{N}}i_{!}^{fake}(\mathcal{E},\alpha)=(\ker(1\otimes\tilde{D}),\,\tilde{T}\otimes 1|_{\ker(1\otimes\tilde{D})}).$
Let $\{s_{i}\}$ be a $\mathbb{C}(\mathcal{Z})$-basis of $C(\mathcal{N})$, and form $\beta_{i},\tilde{D}_{i}$ from the $s_{i}$ as we form $\beta,\tilde{D}$. Then $1\otimes\tilde{D}_{i}$ commutes with $\tilde{T}\otimes 1$ and with $\phi$. If $(V,\tilde{T}\otimes 1)\in K_{0}^{fake}(\mathcal{Z})$, then the self-adjoint operator $\tilde{T}\otimes 1$ almost commutes with the $\mathbb{C}(\mathcal{Z})$-action, and so represents the same element as $(V,\tilde{T}\otimes 1-\varepsilon\tilde{D}_{i})$. Thus, again by the previous lemma, we can restrict $V$ to $\ker\tilde{D}_{i}$. Hence,
$\pi_{\ast}\mu_{\mathcal{N}}i_{!}^{fake}(\mathcal{E},\alpha)=\Bigl{(}\bigcap_{i}\ker(1\otimes\tilde{D}_{i}),\tilde{T}\otimes 1\Bigr{)}.$
On each fiber of $\mathcal{N}\to\mathcal{Z}$, the equations given by $\tilde{D}_{i}=0$ precisely determine a section over $\mathcal{N}$ from its initial values in $\mathcal{Z}$ ([4]). Thus the common kernel of the $\tilde{D}_{i}$ is isomorphic to $(L^{2}(\mathcal{Z},\mathcal{E}),\tilde{T}\otimes 1)=(L^{2}(\mathcal{E}),T\otimes 1)$, and the proposition follows. ∎
@@
Note that the statement on twisted sectors is identical to that on the non-twisted sector (with $\mathcal{Z}$, $\mathcal{N}$, and $\mathcal{Y}$ replaced by the corresponding twisted sectors in these spaces). It suffices to prove the statement sector-wise.
@@
Observe that the contraction $\pi:\mathcal{N}\to\mathcal{Z}$ composed with $i^{fake}:\mathcal{Z}\hookrightarrow\mathcal{N}$ is homotopic to $id:\mathcal{N}\to\mathcal{N}$, thus inducing a homotopy $\pi^{\ast}(i^{fake})^{\ast}\cong id$ on $\mathbb{C}(\mathcal{N})$. By [10] Theorem 3 in Section 3, we know that $i^{fake}_{\ast}\pi_{\ast}=id$ on $K_{0}^{fake}(\mathcal{N})$. On the other hand, it follows from $\pi i^{fake}=id_{\mathcal{Z}}$ that $\pi_{\ast}i^{fake}_{\ast}=id$ on $K_{0}^{fake}(\mathcal{Z})$. So $\pi_{\ast}$ is the inverse of $i^{fake}_{\ast}$, and the corollary follows. ∎
@@
Part (1) follows from the definition. If furthermore $\mathcal{Y}$ deformation retracts to $\mathcal{N}$, then $j_{\ast}$ is an isomorphism by [10] Theorem 3 in Section 3. Thus, $j_{\ast}=(j^{!})^{-1}$ by part (1). ∎
@@
The following diagrams commute for each open embedding $\mathcal{N}\hookrightarrow\mathcal{Y}$.
@@
The commutativity of the first diagram follows from Corollary 2.2. The commutativity of the second diagram follows from that of the first diagram and the fact that $j^{!}j_{\ast}$ is the identity map. ∎
Proposition 3.3 follows from Lemmas 3.10, 3.11, and Corollary 3.1.
@@ 3.5. Push-forward from inertia orbifold to point
Consider sequences
$(g,\,g_{1}G_{1},\,\cdots,\,g_{n}G_{n})$
of cosets in $G$, where $G_{i+1}$ is the group generated by the previous coset $g_{i}G_{i}$ and where $G_{1}$ is the group generated by $g$. Let $\mathcal{G}$ denote the set of such sequences in which, for each $i$, the element $g_{i}$ is not contained in $G_{i}$. Given a sequence $\mathbf{g}=(g,g_{1},\cdots,g_{n})$, we construct a sequence of cosets by letting $G_{i}$ be the group generated by $g,g_{1},\cdots,g_{i}$. In other words, the sequence $\mathbf{g}$ defines an element in $\mathcal{G}$ if and only if each $g_{i}$ is not contained in the group generated by the preceding elements.
Given $\mathbf{g}$, let $\mathcal{Y}^{\mathbf{g}}$ be the $(n+1)$-multi-sector corresponding to $\mathbf{g}$, and let
$i_{\mathbf{g}}:\mathcal{Y}^{\mathbf{g}}\to\mathcal{Y}$
denote the inclusion. Let $n=|\mathbf{g}|$ and let $m(\mathbf{g})$ denote the order of the isotropy group of $\mathcal{Y}^{\mathbf{g}}$. For each $i$, let $\mathcal{N}_{g_{i}}$ be the normal bundle of $\mathcal{Y}^{(g,g_{1},\cdots,g_{i})}$ in $\mathcal{Y}^{(g=g_{0},g_{1},\cdots,g_{i-1})}$, and in particular, let $\mathcal{N}_{g}$ be the normal bundle of $\mathcal{Y}^{(g)}$ in $\mathcal{Y}$. Define
$e^{-1}_{\mathbf{g}}=\Bigl{(}\prod_{i=0}^{n}str_{g_{i}}\,\wedge^{\ast}\Bigl{(}\mathcal{N}^{\ast}_{g_{i}}\otimes\mathbb{C}\Bigr{)}\Bigr{)}^{-1}.$
These objects depend only on the element in $\mathcal{G}$ represented by $\mathbf{g}$ and are independent of the particular representative $\mathbf{g}$.
We define the fake push-forward $p_{\ast}^{fake}$ of an element $(H,\phi,T)$ in the fake $K$-homology to be the Fredholm index of the operator
$T_{+}:H_{+}\to H_{-},$
where $H_{+}$ and $H_{-}$ denote the positive and negative parts of $H$ under the $\mathbb{Z}_{2}$-grading, and where $T_{+}=T|_{H_{+}}$.
@@
The orbifold “push-forward to point” $p_{\ast}(\varphi)$ of
$\{\varphi^{(g)}=(H^{(g)},\phi^{(g)},T^{(g)})\}\in K_{0}(\mathcal{Y})$
is defined by
$p_{\ast}(\varphi)=p_{\ast}^{fake}\!\left(\sum_{\mathbf{g}\in\mathcal{G}}\frac{(-1)^{|\mathbf{g}|+\dim\mathcal{Y}^{(g)}}}{m(\mathbf{g})}\Bigl{(}e_{\mathbf{g}}^{-1}\,i_{\mathbf{g}}^{!}\varphi^{(g)}\Bigr{)}\right).$
Since points in the compact orbifold $\mathcal{Y}$ have finite isotropy, there is a uniform bound $N$ on the orders of the isotropy groups, and $n$-sectors $\mathcal{Y}^{\mathbf{g}}$ become empty when $n>N$. Thus, this is a finite sum.
@@
The following diagram commutes, where $\mathbb{Z}\to\mathbb{C}$ is the natural inclusion and $\operatorname{ind}(\cdots)_{+}$ maps an operator $u$ to the index of its restriction $u_{+}$.